Georgia, Inc. has collected the following data on one of its products. The actual cost of the direct materials used is: Direct materials standard (5 lbs @ $1/lb) $ 5 per finished unit Total direct materials cost variance—unfavorable $ 15,750 Actual direct materials used 150,000 lbs Actual finished units produced 25,000 units

$125,000.

$109,250.

$194,500.

$150,000.

$140,750.
Answer #1

Materials cost variance = Standard cost - Actual cost

(5*25,000) - Actual cost = - 15,750

125,000 - Actual cost = - 15,750

Actual cost = 125,000+15,750 = 140,750

Option E